Find the dual basis of {(1,0,1),(1,1,0),(0,1,1)} in R^3


1
Expert's answer
2021-03-08T19:15:16-0500

We need to find vectors:


"e'_1 = (x_1,y_1,z_1) \\
e'_2 = (x_2,y_2,z_2) \\
e'_3 = (x_3,y_3,z_3) \\"

The condition equals to three systems of three equations. Each system will give us one vector for the dual base.

First vector:

"\begin{cases}
 x_1+z_1=1 \\
 x_1+y_1=0 \\
 y_1+z_1=0 \\
\end{cases}"

"x_1=\frac{1}{2}; y_1=-\frac{1}{2}; z_1=\frac{1}{2}"

"e'_1 = (\frac{1}{2},-\frac{1}{2},\frac{1}{2}) \\"

Second vector:

"\begin{cases}
 x_2+z_2=0 \\
 x_2+y_2=1 \\
 y_2+z_2=0 \\
\end{cases}"

"x_2=\frac{1}{2}; y_2=\frac{1}{2}; z_2=-\frac{1}{2}"

"e'_2 = (\frac{1}{2},\frac{1}{2},-\frac{1}{2}) \\"

Third vector:

"\begin{cases}
 x_3+z_3=0 \\
 x_3+y_3=0 \\
 y_3+z_3=1 \\
\end{cases}"

"x_3=-\frac{1}{2}; y_3=\frac{1}{2}; z_3=\frac{1}{2}"

"e'_3 = (-\frac{1}{2},\frac{1}{2},\frac{1}{2}) \\"

Answer: "e'_1 = (\frac{1}{2},-\frac{1}{2},\frac{1}{2}) \\" ; "e'_2 = (\frac{1}{2},\frac{1}{2},-\frac{1}{2}) \\" ; "e'_3 = (-\frac{1}{2},\frac{1}{2},\frac{1}{2}) \\".